OpenStudy (anonymous):

What is the value of f[g(4)] for the functions f(x) = 2x + 1 and g(x) = 2x - 5? (Type only the value of your final answer, no variables.)

OpenStudy (anonymous):

7

OpenStudy (anonymous):

8-5 then 6 +1

OpenStudy (anonymous):

f(g(x))=2(2x-5)+1 so f(g(4))=2(8-5)+1 =6+1=7